If the marginal propensity to consume is 80 percent, an increase in the national
income of R100 billion would
1
Expert's answer
2021-10-15T10:18:54-0400

Marginal propensity to consume = change in consumption / Change in disposable income

so Marginal propensity to save = 1 - MPC

MPS = 1- MPC

MPS = 1-0.80

MPS = 0.20

0.20 = change in saving / R100 billion

change in saving = R20 billion

saving will be increased by R20 billion.